# Distributional and mean Li–Yorke chaos for weighted shifts on Fréchet sequence spaces

### 02 Mar 2026 v1

@@ 1 Introduction
In [17], sufficient conditions are given under which weighted backward shifts on Köthe sequence spaces $\lambda_{p}(A,\mathbb{N})$ ($p\in[1,\infty)$) are distributionally chaotic. In the first part of this work, our aim is to provide a complete characterization of distributional chaos for weighted backward shifts in a setting more general than Köthe sequence spaces, namely, Fréchet sequence spaces. These spaces are subspaces of $\mathbb{K}^{\mathbb{N}}$ endowed with a topology that turns them into Fréchet spaces and ensures the continuity of the canonical projections. Our approach to obtaining such a characterization is based on the method used in [9], based on the Distributional Chaos Criterion.
@@
The paper is organized as follows. In Section 2, we recall some definitions of Fréchet sequence spaces and fix the notation. In Section 3, we establish a characterization of distributional chaos for weighted backward shifts in the more general setting of Fréchet sequence spaces; as a consequence, we obtain corollaries characterizing this property for Köthe sequence spaces and for the classical Banach spaces $\ell^{p}(\mathbb{Z})$ ($p\in[1,\infty)$) and $c_{0}(\mathbb{Z})$. Finally, in Section 4, we characterize mean Li–Yorke chaos for weighted backward shifts on Fréchet sequence spaces satisfying condition (C), and we derive corresponding corollaries for the same particular cases mentioned above.
@@ 2 Preliminaries
is a continuous linear operator on $X$ provided that it maps $X$ into itself.
@@
is a continuous linear operator on $X$ provided that it maps $X$ into itself.
@@
    $\|x\|_{k}:=\left(\sum_{j\in J}|a_{j,k}x_{j}|^{p}\right)^{1/p},\qquad x=(x_{j})_{j\in J}\in\lambda_{p}(A,J),\;k\in\mathbb{N}.$
@@
    $\|x\|_{k}:=\sup_{j\in J}|a_{j,k}x_{j}|,\qquad x=(x_{j})_{j\in J}\in\lambda_{0}(A,J),\;k\in\mathbb{N}.$
@@
Let $J=\mathbb{N}$ or $J=\mathbb{Z}$ and let $\nu:=(v_{n})_{n\in J}\subset[0,\infty)$. Consider the Köthe matrix $A=(a_{j,k})_{j\in J,\,k\in\mathbb{N}}$ defined by $a_{j,k}=v_{j}$ for all $k\in\mathbb{N}$.
@@
    $c_{0}(\nu,J)=\{(x_{j})_{j\in J}\in\mathbb{K}^{J}:v_{j}x_{j}\to 0\text{ as }|j|\to\infty\}.$
@@
    $\ell^{p}(\nu,J)=\{(x_{j})_{j\in J}\in\mathbb{K}^{J}:\sum_{j\in J}|v_{j}x_{j}|^{p}<\infty\}.$
@@ 3 Distributional chaos
    There exist $m\in\mathbb{N}$ and an increasing sequence $(N_{k})_{k\in\mathbb{N}}$ of positive integers, such that for each $k\in\mathbb{N}$ there are $r:=r(k)\in\mathbb{N}$, indices $i_{1,k},\cdots,i_{r,k}\in I$ and scalars $b_{1,k},\cdots,b_{r,k}\in\mathbb{K}\setminus\left\{0\right\}$ with $\left\|\sum_{j=1}^{r}b_{j,k}e_{i_{j,k}}\right\|_{p(k)}\neq 0$ and
@@
$(\Rightarrow):$ Since $B_{w}$ is distributionally chaotic, by Theorem 11, $B_{w}$ admits a distributionally irregular vector $x:=(x_{n})_{n\in\mathbb{Z}}$, that is, there exist sets $D,E\subset\mathbb{N}$ with $\overline{\text{dens}}(D)=\overline{\text{dens}}(E)=1$ and $m\in\mathbb{N}$ such that
@@
It is easy to see that we can take $m\in\mathbb{N}$ in (2) sufficiently large such that $\left\|x\right\|_{m}\neq 0$. Take $I:=\left\{i\in\mathbb{Z}:x_{i}\neq 0\right\}$. Let $V$ be a neighborhood of 0 in $X$. Then, by the equicontinuity of the family of maps $y:=(y_{n})_{n\in\mathbb{Z}}\in X\mapsto y_{k}e_{k}$ ($k\in\mathbb{Z}$), there exists a neighborhood $U$ of 0 in $X$ such that
@@
$n\in D\,\,\,\wedge\,\,\,n\geq n_{0}\quad\Rightarrow\quad(B_{w})^{n}(x)\in U\quad\Rightarrow\quad x_{i}w_{i-n}\cdots w_{i-1}e_{i-n}\in V,$
@@
$\overline{\text{dens}}\left\{n\in\mathbb{N}:\left\|(B_{w})^{n}(x)\right\|_{m}>k(\left\|x\right\|_{k}+1)\right\}=\overline{\text{dens}}(E)=1,$
@@
($\Leftarrow$): For this implication, we will use the Distributional Chaos Criterion. By item (A), for each $i\in I$ we have that
@@
Let $X$ be a Fréchet sequence, endowed with an increasing sequence $(\left\|\cdot\right\|_{n})_{n\in\mathbb{N}}$ of seminorms, in which the sequence $(e_{n})_{n\in\mathbb{N}}$ of canonical vectors is a basis. Suppose that the unilateral weighted backward shift $B_{w}$, with nonzero weights $w:=(w_{n})_{n\in\mathbb{N}}$, is well-defined and continuous on $X$. Then, $B_{w}$ is distributionally chaotic if and only if the following condition holds:
@@
As a first application, we prove that Theorem 14 recovers [17, Theorem 11], which establishes a sufficient condition in the case of $\ell^{p}(\nu,\mathbb{N})$ for $p\in[1,\infty)$ or $c_{0}(\nu,\mathbb{N})$. To this end, we introduce the following notation:
$S_{i,j}(\alpha)=\{k\in[i,j-1]:a_{k}\geq\alpha\},$
@@
Let $(\nu_{n})_{n\in\mathbb{N}}\subset(0,\infty)$ and fix $p\in\{0\}\cup[1,\infty)$. Assume that the backward shift $B$ is well-defined and continuous on $\ell^{p}(\nu,\mathbb{N})$ (or $c_{0}(\nu,\mathbb{N})$, if $p=0$). If there exist a sequence $(\alpha_{n})_{n\in\mathbb{N}}\subset(0,+\infty)$ and increasing functions $j_{0},j_{1}:\mathbb{N}\to\mathbb{N}$ such that $j_{1}(n)-j_{0}(n)\geq n$ for all
@@
    $\lim_{n\to\infty}\frac{\text{card}\left(S_{j_{0}(n),\,j_{1}(n)}(\alpha_{n})\right)}{j_{1}(n)-j_{0}(n)}=1,$
@@
Without loss of generality, we assume that the sequence $(j_{1}(n)-j_{0}(n))_{n\in\mathbb{N}}$ is strictly increasing. For each $k\in\mathbb{N}$ take $n_{k}\in\mathbb{N}$ such that $\frac{a_{j_{1}(n_{k})}}{\alpha_{n_{k}}}<\frac{1}{2k}$ and $\frac{\text{card}\left(S_{j_{0}(n_{k}),\,j_{1}(n_{k})}(\alpha_{n_{k}})\right)}{j_{1}(n_{k})-j_{0}(n_{k})}>(1-k^{-1}).$ Define $N_{k}:=j_{1}(n_{k})-j_{0}(n_{k})$, $k\in\mathbb{N}$. Then
$\text{card}\left\{1\leq i\leq N_{k}:\frac{\left\|\frac{e_{j_{1}(n_{k})-i}}{\alpha_{n_{k}}}\right\|}{\left\|\frac{e_{j_{1}(n_{k})}}{\alpha_{n_{k}}}\right\|}>k\right\}\geq\text{card}\left(S_{j_{0}(n_{k}),\,j_{1}(n_{k})}(\alpha_{n_{k}})\right)>(1-k^{-1})N_{k}.$
@@
$(w_{n})_{n\in\mathbb{Z}}:=(\cdots,\underset{\text{Block }B_{n}}{\underbrace{\overset{n\text{ times}}{\overbrace{2^{(-1)^{n}},\cdots,2^{(-1)^{n}}}},\overset{n\text{ times}}{\overbrace{2^{(-1)^{n+1}},\cdots,2^{(-1)^{n+1}}}}}},\cdots,\underset{\text{Block }B_{2}}{\underbrace{2,2,\frac{1}{2},\frac{1}{2}}},\underset{\text{Block }B_{1}}{\underbrace{\frac{1}{2},2}},\underset{j\geq 0}{\underbrace{2,2,2,2,\cdots}}).$
For each $n\in\mathbb{N}$, denote by $I_{n}$ the set of indices that compose the Block $n$. Moreover, we denote $I_{n}:=\left\{-j:j\in I_{n}\right\}$ ($n\in\mathbb{N}$). We need the following lemma to prove that $B_{w}$ is not distributionally chaotic:
@@
The literature already contains examples of distributionally chaotic shifts that are not hypercyclic; see, for instance, [17, Theorem 11]. In what follows, in order to illustrate the use of the sufficiency direction in Theorem 12 (in particular Corollary 16), we present an example of a weighted backward shift on $\lambda_{p}(A,\mathbb{Z})$ ($p\in\{0\}\cup[1,\infty)$), which is distributionally chaotic but not hypercyclic.
@@
We denote by $I_{n}$ the set of indices that compose the block $B_{n,k}$ $(n,k\in\mathbb{N})$, note that these indices do not depend on $k$. Now, consider the weighted backward shift $B_{w}$ on $\lambda_{p}(A,\mathbb{Z})$ ($p\in\left\{0\right\}\cup[1,\infty)$) with weights $w_{n}:=\frac{1}{2}$, for $n<0$ and $w_{n}:=1$, for $n\geq 0$. Since
@@
We use the following lemma in the next corollary.
Lemma 20.
Consider $X:=\ell^{p}(\mathbb{Z})$ ($1\leq p<\infty$) or $X:=c_{0}(\mathbb{Z})$. Let $w:=(w_{n})_{n\in\mathbb{N}}$ be a sequence of nonzero scalars such that the bilateral weighted backward shift $B_{w}$ is a well-defined and continuous operator on $X$. Then, the following conditions are equivalent:
  • (a)
    There exist $\varepsilon>0$ and an increasing sequence $(N_{k})_{k\in\mathbb{N}}$ of positive integers such that, for each $k\in\mathbb{N}$, there are $r\in\mathbb{N}$, indices $i_{1},\ldots,i_{r}\in\mathbb{N}$ and scalars $b_{1},\ldots,b_{r}\in(0,\infty)$ with
    $\operatorname{card}\left\{1\leq n\leq N_{k}:\frac{\left\|(B_{w})^{n}\left(\sum_{j=1}^{r}b_{j}e_{i_{j}}\right)\right\|}{\left\|\sum_{j=1}^{r}b_{j}e_{i_{j}}\right\|}>k\right\}>\varepsilon N_{k}.$
  • (b)
    There exist an increasing sequence $(N^{\prime}_{k})_{k\in\mathbb{N}}$ of positive integers such that, for each $k\in\mathbb{N}$, there are $r^{\prime}\in\mathbb{N}$, indices $i^{\prime}_{1},\ldots,i^{\prime}_{r^{\prime}}\in\mathbb{N}$ and scalars $b^{\prime}_{1},\ldots,b^{\prime}_{r^{\prime}}\in(0,\infty)$ with
    $\operatorname{card}\left\{1\leq n\leq N^{\prime}_{k}:\frac{\left\|(B_{w})^{n}\left(\sum_{j=1}^{r^{\prime}}b^{\prime}_{j}e_{i^{\prime}_{j}}\right)\right\|}{\left\|\sum_{j=1}^{r^{\prime}}b^{\prime}_{j}e_{i^{\prime}_{j}}\right\|}>k\right\}>(1-k^{-1})N^{\prime}_{k}.$
Proof.
$((b)\Rightarrow(a)):$ It is trivial.
$((a)\Rightarrow(b)):$ By [5, Proposition 8], there exist $\varepsilon>0$, a sequence $(y_{k})$ in $X$ and an increasing sequence $(N_{k})$ in $\mathbb{N}$ such that $\lim_{k\to\infty}y_{k}=0$ and
$\lim_{k\to\infty}\frac{1}{N_{k}}\operatorname{card}\{1\leq j\leq N_{k}:\left\|(B_{w})^{n}(y_{k})\right\|>\varepsilon\}=1.$
Consider a subsequence $(y_{k_{s}})_{s\in\mathbb{N}}$ such that $(k_{s})_{s\in\mathbb{N}}$ is an increasing sequence,
$\left\|y_{k_{s}}\right\|<\frac{\varepsilon}{2s}\quad\text{and}\quad\operatorname{card}\{1\leq n\leq N_{k_{s}}:\left\|(B_{w})^{n}(y_{k_{s}})\right\|>\varepsilon\}>(1-(2s)^{-1})N_{k_{s}},$ (13)
for all $s\in\mathbb{N}$. Then, (13) implies that
$\operatorname{card}\left\{1\leq n\leq N_{k_{s}}:\frac{\left\|(B_{w})^{n}(y_{k_{s}})\right\|}{\left\|y_{k_{s}}\right\|}>2s\right\}>(1-(2s)^{-1})N_{k_{s}}.$ (14)
By the density of $c_{00}(\mathbb{Z})$ in $X$, for each $s\in\mathbb{N}$ there exist $r^{\prime}\in\mathbb{N}$, indices $i^{\prime}_{1},\ldots,i^{\prime}_{r^{\prime}}\in\mathbb{Z}$ and scalars $b^{\prime}_{1},\ldots,b^{\prime}_{r^{\prime}}\in(0,\infty)$ such that
$\left|\frac{\left\|(B_{w})^{n}(y_{k_{s}})\right\|}{\left\|y_{k_{s}}\right\|}-\frac{\left\|(B_{w})^{n}\left(\sum_{j=1}^{r^{\prime}}b^{\prime}_{j}e_{i^{\prime}_{j}}\right)\right\|}{\left\|\sum_{j=1}^{r^{\prime}}b^{\prime}_{j}e_{i^{\prime}_{j}}\right\|}\right|<1,\quad\text{ for all }1\leq n\leq N_{k_{s}}.$ (15)
Now, put $N^{\prime}_{s}:=N_{k_{s}}$, for each $s\in\mathbb{N}$. Then $(N^{\prime}_{s})_{s\in\mathbb{N}}$ is an increasing sequence. By (14) and (15), we obtain
$\operatorname{card}\left\{1\leq n\leq N^{\prime}_{s}:\frac{\left\|(B_{w})^{n}\left(\sum_{j=1}^{r^{\prime}}b^{\prime}_{j}e_{i^{\prime}_{j}}\right)\right\|}{\left\|\sum_{j=1}^{r^{\prime}}b^{\prime}_{j}e_{i^{\prime}_{j}}\right\|}>s\right\}>(1-s^{-1})N^{\prime}_{s}.$
In the following corollary, we show that Corollary 16 extends the characterizations of distributional chaos for weighted backward shifts on the spaces $\ell^{p}(\mathbb{Z})$ and $c_{0}(\mathbb{Z})$, given in [9].
Corollary 21.
Consider $X:=\ell^{p}(\mathbb{Z})$ ($1\leq p<\infty$) or $X:=c_{0}(\mathbb{Z})$. Let $w:=(w_{n})_{n\in\mathbb{N}}$ be a sequence of nonzero scalars such that the bilateral weighted backward shift $B_{w}$ is a well-defined and continuous operator on $X$.
  • (a)
    If $X=c_{0}(\mathbb{Z})$, then $B_{w}$ is distributionally chaotic on $c_{0}(\mathbb{Z})$ if and only if there exists a subset $S\subset\mathbb{Z}$ such that the following conditions are satisfied:
    • (A1)
      There exists $D\subset\mathbb{N}$ with $\operatorname{\overline{dens}}(D)=1$ and $\lim_{n\in D}|w_{i-n}\cdots w_{i-1}|=0$, for all $i\in S$.
    • (A2)
      $\inf_{k\in\mathbb{N}}\left(\sup_{N\in\mathbb{N}}\frac{\operatorname{card}\{1\leq n\leq N:|w_{i-n}\cdots w_{i-1}|>k\text{ for some }i\in S\}}{N}\right)>0.$
  • (b)
    If $X=\ell^{p}(\mathbb{Z})$, then $B_{w}$ is distributionally chaotic on $\ell^{p}(\mathbb{Z})$ if and only if there exists a subset $S\subset\mathbb{Z}$ such that the following conditions are satisfied:
    • (B1)
      There exists $D\subset\mathbb{N}$ with $\operatorname{\overline{dens}}(D)=1$ and $\lim_{n\in D}|w_{i-n}\cdots w_{i-1}|=0$, for all $i\in S$.
    • (B2)
      There exist $\varepsilon>0$ and an increasing sequence $(N_{k})_{k\in\mathbb{N}}$ of positive integers such that, for each $k\in\mathbb{N}$, there are $r\in\mathbb{N}$, indices $i_{1},\dots,i_{r}\in S$ and scalars $b_{1},\dots,b_{r}\in(0,\infty)$ with
      $\operatorname{card}\left\{1\leq n\leq N_{k}:\frac{b_{1}|w_{i_{1}-n}\cdots w_{i_{1}-1}|^{p}\;+\;\cdots\;+\;b_{r}|w_{i_{r}-n}\cdots w_{i_{r}-1}|^{p}}{b_{1}+\cdots+b_{r}}>k\right\}\;\geq\;\varepsilon N_{k}.$ (16)
Proof.
(a) First, recall that we can view the space $c_{0}(\mathbb{Z})$ as the space $\lambda_{0}(A,\mathbb{Z})$ where $A=(a_{j,k})_{j\in\mathbb{Z},k\in\mathbb{N}}$ is defined as $a_{j,k}=1$ for all $j\in\mathbb{Z}$ and $k\in\mathbb{N}$. Then, the condition (A1) is equivalent to the condition (A1) of Corollary 16. Now, we will prove that condition (A2) is equivalent to the condition (A2) of Corollary 16.
Suppose that condition (A2) of the statement holds. Then, there are $\varepsilon>0$ and an increasing sequence $(N_{k})_{k\in\mathbb{N}}$ of positive integers, such that
$\frac{\operatorname{card}\{1\leq n\leq N_{k}:|w_{i-n}\cdots w_{i-1}|>k\text{ for some }i\in S\}}{N_{k}}>\varepsilon,\quad\text{for each }k\in\mathbb{N}.$ (17)
Define $I_{k}:=\{\,1\leq n\leq N_{k}:|w_{i-n}\cdots w_{i-1}|>k\text{ for some }i\in S\,\}.$ For each $n\in I_{k}$, take $i_{k,n}\in S$ such that $|w_{i_{k,n}-n}\cdots w_{i_{k,n}-1}|>k.$ Then, by (17), we have
$\text{card}\left\{1\leq n\leq N_{k}:\left\|(B_{w})^{n}\left(\sum_{j\in I_{k}}e_{i_{k,j}}\right)\right\|>k\right\}>N_{k}\varepsilon.$
Thus, by Lemma 20, we obtain the condition (A2) of Corollary 16. Now, suppose that the condition (A2) of Corollary 16 holds. Since,
$\max_{1\leq j\leq r}|w_{i_{j,k}-n}\cdots w_{i_{j,k}-1}|\geq\frac{\max_{1\leq j\leq r}|b_{j,k}w_{i_{j,k}-n}\cdots w_{i_{j,k}-1}|}{\max_{1\leq j\leq r}\left|b_{j,k}\right|},$
then, (6) implies that
$\text{card}\left\{1\leq n\leq N_{k}:\max_{1\leq j\leq r}|w_{i_{j,k}-n}\cdots w_{i_{j,k}-1}|>k\right\}>(1-k^{-1})N_{k},$
for all $k\in\mathbb{N}$. Then, there exist $\varepsilon>0$ such that
$\sup_{N\in\mathbb{N}}\frac{1}{N}\text{card}\left\{1\leq n\leq N:|w_{i-n}\cdots w_{i-1}|>k,\,\text{for some }i\in\mathbb{Z}\right\}>\varepsilon,\quad\text{for all }k\in\mathbb{N}.$
Therefore, the condition (A2) of our enunciate holds.
(b) First, for a fixed $p\in[1,\infty)$, recall that we can view the space $\ell^{p}(\mathbb{Z})$ as the space $\lambda_{p}(A,\mathbb{Z})$ where $A=(a_{j,k})_{j\in\mathbb{Z},k\in\mathbb{N}}$ is defined as $a_{j,k}=1$ for all $j\in\mathbb{Z}$ and $k\in\mathbb{N}$. Then, the condition (B1) is equivalent to the condition (B1) of Corollary 16. Now, note that (16) is equivalent to
$\operatorname{card}\left\{1\leq n\leq N_{k}:\frac{\left\|(B_{w})^{n}\left(\sum_{j=1}^{r}b_{j}^{\frac{1}{p}}e_{j}\right)\right\|}{\left\|\sum_{j=1}^{r}b_{j}^{\frac{1}{p}}e_{j}\right\|}>k^{\frac{1}{p}}\right\}\;\geq\;\varepsilon N_{k},\quad\text{for all }k\in\mathbb{N}.$
Then, by Lemma 20, this is equivalent to the following condition: there exists an increasing sequence $(N^{\prime}_{k})_{k\in\mathbb{N}}$ of positive integers such that, for each $k\in\mathbb{N}$, there are $r^{\prime}\in\mathbb{N}$, indices $i^{\prime}_{1},\ldots,i^{\prime}_{r}\in S$ and scalars $b^{\prime}_{1},\ldots,b^{\prime}_{r}\in(0,\infty)$ with
$\operatorname{card}\left\{1\leq n\leq N^{\prime}_{k}:\frac{\left\|(B_{w})^{n}\left(\sum_{j=1}^{r^{\prime}}b^{\prime\frac{1}{p}}_{j}e_{i^{\prime}_{j}}\right)\right\|}{\left\|\sum_{j=1}^{r^{\prime}}b^{\prime\frac{1}{p}}_{j}e_{i^{\prime}_{j}}\right\|}>k\right\}>(1-k^{-1})N^{\prime}_{k}.$
The latter condition is equivalent to the condition (B2) of Corollary 16. ∎
Remark 22.
Using Theorem 14, one obtains unilateral versions of Corollaries 16 and 21 by arguments analogous to the corresponding proofs. We leave the details to the reader.

### 06 Apr 2026 v2

@@ 1 Introduction
In [18], sufficient conditions are given under which backward shifts on Köthe sequence spaces $\lambda_{p}(A,\mathbb{N})$ ($p\in[1,\infty)\cup\{0\}$) are distributionally chaotic. In the first part of this work, our aim is to provide a complete characterization of distributional chaos for weighted backward shifts in a setting more general than Köthe sequence spaces, namely, Fréchet sequence spaces. These spaces are subspaces of $\mathbb{K}^{\mathbb{N}}$ endowed with a topology that turns them into Fréchet spaces and ensures the continuity of the canonical projections. Our approach to obtaining such a characterization is based on the method used in [9], based on the Distributional Chaos Criterion.
@@
In order to mention other related works, we refer to [17, 20]. In [17], the authors study distributional chaos for weighted backward shifts on spaces of the form $\Sigma(X):=X^{\mathbb{N}}$, where $X$ is a Banach space. In [20], the authors provide a characterization of Li–Yorke chaos for (unweighted) shifts on Köthe sequence spaces $\lambda_{p}(A,\mathbb{N})$, with $p\in[1,\infty]\cup\{0\}$.
@@
The paper is organized as follows. In Section 2, we recall some definitions of Fréchet sequence spaces and fix the notation. In Section 3, we establish a characterization of distributional chaos for weighted backward shifts in the more general setting of Fréchet sequence spaces; as a consequence, we obtain corollaries characterizing this property for Köthe sequence spaces. Finally, in Section 4, we characterize mean Li–Yorke chaos for weighted backward shifts on Fréchet sequence spaces satisfying condition (C).
@@ 2 Preliminaries
is a continuous linear operator on $X$ provided that it maps $X$ into itself. If $w:=(1)_{n\in\mathbb{N}}$, then we denote $B_{w}=B$.
@@
is a continuous linear operator on $X$ provided that it maps $X$ into itself. If $w:=(1)_{n\in\mathbb{Z}}$, then we denote $B_{w}=B$.
@@
    $\|x\|_{k}:=\left(\sum_{j\in J}|a_{j,k}x_{j}|^{p}\right)^{1/p},\quad\text{ for }\;x=(x_{j})_{j\in J}\in\lambda_{p}(A,J)\;\text{ and }k\in\mathbb{N}.$
@@
    $\|x\|_{k}:=\sup_{j\in J}|a_{j,k}x_{j}|,\quad\text{ for }\;x=(x_{j})_{j\in J}\in\lambda_{0}(A,J)\;\text{ and }k\in\mathbb{N}.$
@@
Let $J=\mathbb{N}$ or $J=\mathbb{Z}$ and let $\nu:=(v_{n})_{n\in J}\subset(0,\infty)$. Consider the Köthe matrix $A=(a_{j,k})_{j\in J,\,k\in\mathbb{N}}$ defined by $a_{j,k}=v_{j}$ for all $k\in\mathbb{N}$.
@@
    $c_{0}(\nu,J)=\{(x_{j})_{j\in J}\in\mathbb{K}^{J}:v_{j}x_{j}\to 0\text{ as }|j|\to\infty\},\quad\text{with norm }\left\|(x_{n})_{n\in J}\right\|=\sup_{n\in J}\left|v_{n}x_{n}\right|.$
@@
    $\ell^{p}(\nu,J)=\{(x_{j})_{j\in J}\in\mathbb{K}^{J}:\sum_{j\in J}|v_{j}x_{j}|^{p}<\infty\}\quad\text{with norm }\left\|(x_{n})_{n\in J}\right\|=\left(\sum_{j\in J}|v_{j}x_{j}|^{p}\right)^{\frac{1}{p}}.$
@@ 3 Distributional chaos
    There exist $m\in\mathbb{N}$ and an increasing sequence $(N_{k})_{k\in\mathbb{N}}$ of positive integers such that for each $k\in\mathbb{N}$ there are $r:=r(k)\in\mathbb{N}$, indices $i_{1,k},\cdots,i_{r,k}\in I$ and scalars $b_{1,k},\cdots,b_{r,k}\in\mathbb{K}\setminus\left\{0\right\}$ with $\left\|\sum_{j=1}^{r}b_{j,k}e_{i_{j,k}}\right\|_{p(k)}\neq 0$ and
@@
$(\Rightarrow)$ Since $B_{w}$ is distributionally chaotic, by Theorem 11, $B_{w}$ admits a distributionally irregular vector $x:=(x_{n})_{n\in\mathbb{Z}}$, that is, there exist sets $D,E\subset\mathbb{N}$ with $\overline{\text{dens}}(D)=\overline{\text{dens}}(E)=1$ and $m\in\mathbb{N}$ such that
@@
It is easy to see that we can take $m\in\mathbb{N}$ in (2) sufficiently large such that $\left\|x\right\|_{m}\neq 0$. Take $I:=\left\{i\in\mathbb{Z}:x_{i}\neq 0\right\}$. Let $V$ be a neighborhood of 0 in $X$. Then, by the equicontinuity of the family of maps $y:=(y_{n})_{n\in\mathbb{Z}}\in X\mapsto y_{k}e_{k}\in X$ ($k\in\mathbb{Z}$), there exists a neighborhood $U$ of 0 in $X$ such that
@@
$n\in D\,\,\,\text{and}\,\,\,n\geq n_{0}\quad\Rightarrow\quad(B_{w})^{n}(x)\in U\quad\Rightarrow\quad x_{i}w_{i-n}\cdots w_{i-1}e_{i-n}\in V,$
@@
$\overline{\text{dens}}\left\{n\in\mathbb{N}:\left\|(B_{w})^{n}(x)\right\|_{m}>k(\left\|x\right\|_{p(k)}+1)\right\}=\overline{\text{dens}}(E)=1,$
@@
$(\Leftarrow)$ For this implication, we will use the Distributional Chaos Criterion. By item (A), for each $i\in I$ we have that
@@
Let $X$ be a Fréchet sequence space, endowed with an increasing sequence $(\left\|\cdot\right\|_{n})_{n\in\mathbb{N}}$ of seminorms, in which the sequence $(e_{n})_{n\in\mathbb{N}}$ of canonical vectors is a basis. Suppose that the unilateral weighted backward shift $B_{w}$, with nonzero weights $w:=(w_{n})_{n\in\mathbb{N}}$, is well-defined and continuous on $X$. Then, $B_{w}$ is distributionally chaotic if and only if the following condition holds:
@@
As a first application, we prove that Theorem 14 recovers [18, Theorem 11], which establishes a sufficient condition in the case of $\ell^{p}(\nu,\mathbb{N})$ for $p\in[1,\infty)$ or $c_{0}(\nu,\mathbb{N})$. To this end, we introduce the following notation:
$S_{i,j}(\alpha)=\{k\in[i,j]\cap\mathbb{N}:a_{k}\geq\alpha\},$
@@
Let $(\nu_{n})_{n\in\mathbb{N}}\subset(0,\infty)$ and fix $p\in\{0\}\cup[1,\infty)$. Assume that the backward shift $B$ is well-defined and continuous on $\ell^{p}(\nu,\mathbb{N})$ (or $c_{0}(\nu,\mathbb{N})$, if $p=0$). If there exist a sequence $(\alpha_{n})_{n\in\mathbb{N}}\subset(0,+\infty)$ and increasing functions $j_{0},j_{1}:\mathbb{N}\to\mathbb{N}$ such that $j_{1}(n)-j_{0}(n)\geq n$ for all $n\in\mathbb{N}$ and
@@
    $\lim_{n\to\infty}\frac{\operatorname{card}\left(S_{j_{0}(n),\,j_{1}(n)}(\alpha_{n})\right)}{j_{1}(n)-j_{0}(n)}=1,$
@@
Without loss of generality, we assume that the sequence $(j_{1}(n)-j_{0}(n))_{n\in\mathbb{N}}$ is strictly increasing. For each $k\in\mathbb{N}$ take $n_{k}\in\mathbb{N}$ such that $\frac{a_{j_{1}(n_{k})}}{\alpha_{n_{k}}}<\frac{1}{2k}$ and $\frac{\text{card}\left(S_{j_{0}(n_{k}),\,j_{1}(n_{k})}(\alpha_{n_{k}})\right)-1}{j_{1}(n_{k})-j_{0}(n_{k})}>(1-k^{-1}).$ Define $N_{k}:=j_{1}(n_{k})-j_{0}(n_{k})$, $k\in\mathbb{N}$. Then
$\text{card}\left\{1\leq i\leq N_{k}:\frac{\left\|\frac{e_{j_{1}(n_{k})-i}}{\alpha_{n_{k}}}\right\|}{\left\|\frac{e_{j_{1}(n_{k})}}{\alpha_{n_{k}}}\right\|}>k\right\}\geq\text{card}\left(S_{j_{0}(n_{k}),\,j_{1}(n_{k})}(\alpha_{n_{k}})\right)-1>(1-k^{-1})N_{k}.$
@@
Remark 17.
For a fixed $p\in[1,\infty)$, recall that $\ell^{p}(\mathbb{Z})=\lambda_{p}(A,\mathbb{Z})$, where $A=(a_{j,k})_{j\in\mathbb{Z},\,k\in\mathbb{N}}$ is given by $a_{j,k}=1$ for all $j\in\mathbb{Z}$ and $k\in\mathbb{N}$. Likewise, $c_{0}(\mathbb{Z})=\lambda_{0}(A,\mathbb{Z})$. Therefore, by Corollary 16, we obtain characterizations of distributional chaos for weighted shifts on $c_{0}(\mathbb{Z})$ and $\ell^{p}(\mathbb{Z})$ equivalent to those given in [9]. A unilateral version of Corollary 16 follows from Theorem 14. We leave the details to the reader.
@@
$(w_{j})_{j\in\mathbb{Z}}:=(\cdots,\underset{\text{Block }B_{n}}{\underbrace{\overset{n\text{ times}}{\overbrace{2^{(-1)^{n}},\cdots,2^{(-1)^{n}}}},\overset{n\text{ times}}{\overbrace{2^{(-1)^{n+1}},\cdots,2^{(-1)^{n+1}}}}}},\cdots,\underset{\text{Block }B_{2}}{\underbrace{2,2,\frac{1}{2},\frac{1}{2}}},\underset{\text{Block }B_{1}}{\underbrace{\frac{1}{2},2}},\underset{j\geq 0}{\underbrace{2,2,2,2,\cdots}}).$
For each $n\in\mathbb{N}$, denote by $I_{n}$ the set of indices that compose the Block $n$. For example, $I_{1}=\{-2,-1\}$, $I_{2}=\{-6,-5,-4,-3\}$, and so on. Moreover, we denote $I_{n}:=\left\{-j:j\in I_{n}\right\}$ ($n\in\mathbb{N}$). We need the following lemma to prove that $B_{w}$ is not distributionally chaotic:
@@
The literature already contains examples of distributionally chaotic shifts that are not hypercyclic; see, for instance, [18, Example 13]. In what follows, in order to illustrate the use of the sufficiency direction in Theorem 12 (in particular Corollary 16), we present an example of a weighted backward shift on $\lambda_{p}(A,\mathbb{Z})$ ($p\in\{0\}\cup[1,\infty)$), which is distributionally chaotic but not hypercyclic.
@@
We denote by $I_{n}$ the set of indices that compose the block $B_{n,k}$ $(n,k\in\mathbb{N})$, note that these indices do not depend on $k$. For example $I_{1}=\left\{2,3,\cdots,11\right\}$. Now, consider the weighted backward shift $B_{w}$ on $\lambda_{p}(A,\mathbb{Z})$ ($p\in\left\{0\right\}\cup[1,\infty)$) with weights $w_{n}:=\frac{1}{2}$, for $n<0$ and $w_{n}:=1$, for $n\geq 0$. Since